OpenStudy (anonymous):

solve (x-5/x )-1=x+3/x

OpenStudy (e.mccormick):

This is all about order of operations. You need to get x alone. Do you know how you would go about that?

OpenStudy (anonymous):

not really

OpenStudy (e.mccormick):

OK. If you look at it, everything on the right has an x. But the left is mixed up. So why not move things over.

OpenStudy (anonymous):

could you add one to both sides

OpenStudy (e.mccormick):

Yes, but that won't get all the xes together.

OpenStudy (e.mccormick):

\(\left(x-\frac{5}{x} \right)-1=x+\frac{3}{x}\) \(\left(x-\frac{5}{x} \right)-\left(x-\frac{5}{x} \right)-1=x+\frac{3}{x}-\left(x-\frac{5}{x} \right)\) \(-1=x+\frac{3}{x}-\left(x-\frac{5}{x} \right)\)

OpenStudy (e.mccormick):

See, that gets all the Xes on one side. Now, it still needs to be simplified.

OpenStudy (anonymous):

-2/x on the right side

OpenStudy (e.mccormick):

Yes! \(-1=-\frac{2}{x}\) So, what do you think will be good to do next?

OpenStudy (anonymous):

could you multiply both sides by -2

OpenStudy (e.mccormick):

Well, that would still leave the x on the bottom of a fraction. The goal here is to get rid of the fraction next.

OpenStudy (anonymous):

divide by-2

OpenStudy (e.mccormick):

Still leaves a fraction... so not the best choice. Great you are thinking about it, but there is a better choice.

OpenStudy (anonymous):

i dont know

OpenStudy (e.mccormick):

Ever heard about multiplying by the inverse to get rid of a fraction?

OpenStudy (anonymous):

oh

OpenStudy (e.mccormick):

We don't need the whole inverse this time. Mostly we just need the x.

OpenStudy (e.mccormick):

In fact... you could finish this in one step if you multiply by a negative x. That would take care of the - on the 1 and the x under the fraction all at the same time! =)

OpenStudy (anonymous):

so it would be x=2

OpenStudy (e.mccormick):

Yep! \(-1\cdot (-x)=-\frac{2}{x}\cdot (-x)\implies x=2\)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wait thats not an answer choice

OpenStudy (e.mccormick):

np. This is all about figuring out how to get the x alone and out of any fractions. That is all about using the order of operation rules. If you work with those more, these should be a lot easier for you. =)

OpenStudy (e.mccormick):

Oops! I missed one mistake!

OpenStudy (anonymous):

8 0 −1 −8

OpenStudy (e.mccormick):

Look at this again: \(-1=x+\frac{3}{x}-\left(x-\frac{5}{x} \right)\) The negative distribution! \(-1=x+\frac{3}{x}-x+\frac{5}{x} \)

OpenStudy (e.mccormick):

\(-1=\frac{8}{x}\)

OpenStudy (anonymous):

isnt it x-5/x

OpenStudy (e.mccormick):

The - outdide the parentisis distributes to both.